Gamma Industries has net income of $700,000, and it has 465,000 shares of common stock outstanding. The company's stock currently trades at $71 a share. Gamma is considering a plan in which it will use available cash to repurchase 30% of its shares in the open market at the current $71 stock price. The repurchase is expected to have no effect on net income or the company's P/E ratio. What will be its stock price following the stock repurchase? Do not round intermediate calculations. Round your answer to the nearest cent.

$
Answer #1

Solution:-

Existing P/E ratio= Existing price/existing EPS

Or,

Existing P/E ratio= Existing price/(Net income/existing no. of shares)= 71/(700,000/465,000)= 71/1.5054= 47.164 times

Number of shares after repurchase= Current shares outstanding*(1-30%)= 465,000*(1-0.3)= 325,500 shares

Earnings per share after share repurchase= Net income/Number of shares after repurchase= 700,000/325,500= $2.1505 per share

Price of share after repurchase= EPS after repurchase*P/E ratio= 2.1505*47.164= $101.43 per share